OpenStudy (anonymous):

I am having a hard time continuing to the next step with this lim h -->0 f(x+h)-f(x)/h when f(X) = ln x

OpenStudy (anonymous):

i know it will be \[\lim_{h \rightarrow 0} (ln (x+h) - lnx) / h\]

OpenStudy (anonymous):

but i do not know the next step after that.

OpenStudy (anonymous):

very difficult i think the next step is use the property of logariths to make ln( (x+h)/x )

OpenStudy (anonymous):

Use a logarithm property. In the numerator, you have \[\ln(x+h)-\ln x=\ln\frac{x+h}{x}=\ln\left(1+\frac{h}{x}\right)\] \[\lim_{h\to0}\frac{\ln\left(1+\frac{h}{x}\right)}{h}\] Let \(n=\dfrac{1}{h}\). As \(h\to0\), we have \(n\to\infty\). \[\lim_{n\to\infty}n\ln\left(1+\frac{1}{nx}\right)\\ \lim_{n\to\infty}\ln\left(1+\frac{1}{nx}\right)^n\] Next, you multiply by \(\dfrac{x}{x}\): \[\lim_{n\to\infty}\frac{x}{x}\ln\left(1+\frac{1}{nx}\right)^n\\ \lim_{n\to\infty}\frac{1}{x}\ln\left(1+\frac{1}{nx}\right)^{nx}\\ \frac{1}{x}\lim_{n\to\infty}\ln\left(1+\frac{1}{nx}\right)^{nx}\] Since the natural logarithm function is continuous for \(n>0\), you can "pull out" the log: \[\frac{1}{x}\ln\left[\lim_{n\to\infty}\left(1+\frac{1}{nx}\right)^{nx}\right]\]

OpenStudy (anonymous):

that demostration has a lot of math tricks well here is a video about that problem http://www.youtube.com/watch?v=ufyWicnng_E

OpenStudy (anonymous):

Thank You so much!